# Advancing Neural Network Performance through Emergence-Promoting Initialization Scheme

### v1

@@ Abstract
We introduce a novel yet straightforward neural network initialization scheme that modifies conventional methods like Xavier and Kaiming initialization. Inspired by the concept of emergence and leveraging the emergence measures proposed by Li (2023), our method adjusts the layer-wise weight scaling factors to achieve higher emergence values. This enhancement is easy to implement, requiring no additional optimization steps for initialization compared to GradInit. We evaluate our approach across various architectures, including MLP and convolutional architectures for image recognition, and transformers for machine translation. We demonstrate substantial improvements in both model accuracy and training speed, with and without batch normalization. The simplicity, theoretical innovation, and demonstrable empirical advantages of our method make it a potent enhancement to neural network initialization practices. These results suggest a promising direction for leveraging emergence to improve neural network training methodologies. Code is available at: https://github.com/johnnyjingzeli/EmergenceInit.
@@ 1 Introduction
Emergence in machine learning refers to the phenomenon where complex behaviors and properties arise from the interactions of simpler elements within a system. This concept, deeply rooted in the study of complex systems, has gained increasing attention in the context of neural networks. The emergent properties are not merely the sum of individual components but rather the result of intricate interactions among them, leading to sophisticated and often unexpected capabilities. These emergent behaviors are crucial for enabling neural networks to perform complex tasks such as image recognition, natural language processing, and strategic game playing.
The concept of emergence is particularly relevant in neural networks, where the interplay of numerous neurons can give rise to high-level abstractions and decision-making capabilities. Li (2023) provides a framework to quantify emergence in neural networks through the measure of paths from inactive to active nodes. This metric offers a novel lens through which we can understand and enhance the emergent properties of neural networks, potentially leading to more efficient and effective models.
In this paper, we introduce a new initialization scheme for neural networks that leverages the concept of emergence. Our method adjusts the layer-wise variance parameters to achieve higher emergence values compared to traditional methods like Xavier and Kaiming initialization. This approach is particularly appealing because it is straightforward to implement, requiring only minor modifications to existing initialization techniques without necessitating additional optimization or training steps.
@@
Emergence in neural networks can be seen as a manifestation of the broader principle that simple rules or interactions can lead to complex and adaptive behaviors. This principle has been explored in various contexts within machine learning, from the self-organizing properties of unsupervised learning algorithms to the adaptive behaviors of reinforcement learning agents. The work by Li (2023) extends this understanding by providing a concrete measure of emergence that can be directly applied to the design and initialization of neural networks.
Our experiments on standard datasets such as ImageNet for image recognition and the WMT news translation task for machine translation underscore the practical benefits of our initialization scheme. The results highlight not only improvements in final model accuracy but also faster convergence during training, suggesting that our method effectively primes the network for learning by optimizing its emergent properties from the outset.
@@
In summary, our proposed initialization scheme represents a novel and practical approach to improving neural network performance by leveraging the concept of emergence. By adjusting layer-wise variance parameters to enhance emergent properties, we provide a simple yet powerful modification to existing initialization methods. The empirical results presented in this paper demonstrate the significant advantages of our approach, offering a promising direction for future research and applications in machine learning.
@@ 2 Related Work
The initialization of neural networks has been a critical area of research, influencing the stability and speed of training, as well as the ultimate performance of the models. Traditional initialization schemes such as Xavier[12] and Kaiming [13] initialization have laid the foundation for effectively training deep networks by mitigating issues related to vanishing and exploding gradients. Xavier initialization aims to keep the scale of the gradients approximately the same in all layers, while Kaiming initialization, specifically designed for ReLU activations, helps to maintain the variance of activations throughout the layers. Both methods have proven to be fundamental in training deep networks but do not explicitly account for emergent properties within the networks.
@@
Recent research has explored more sophisticated initialization strategies that leverage the structural and statistical properties of neural networks. For instance, Saxe et al. (2013) studied the dynamics of signal propagation in deep networks, highlighting the importance of properly scaling the initial weights to ensure efficient training. Additionally, Mishkin and Matas (2015) proposed a layer-sequential unit-variance (LSUV) initialization that iteratively adjusts the weights to achieve unit variance across all layers, further improving convergence.
@@
The concept of emergence, where complex behaviors arise from simple interactions within a system, has also been examined in the context of neural networks. Emergent properties have been shown to play a crucial role in the development of robust and adaptive models. Research by Olah et al. (2020) has illustrated how higher-level features and behaviors emerge in deep networks as a result of training on large datasets. This phenomenon underscores the potential for leveraging emergent properties to enhance network performance.
(Li, 2023) introduced a novel framework for measuring emergence in neural networks, focusing on the number of paths from inactive to active nodes as a key metric. This work provides a theoretical basis for understanding how emergent behaviors can be quantified and optimized, offering new insights into the design and initialization of neural networks.
Our proposed initialization scheme builds on these foundational studies by explicitly targeting the enhancement of emergent properties from the outset. By adjusting layer-wise variance parameters to achieve higher emergence values, our method represents a simple yet effective modification to traditional initialization techniques. This approach aligns with the growing body of research that seeks to integrate principles from complex systems and emergence theory into machine learning practices.
@@
Furthermore, our work is related to studies that explore the role of initialization in specific architectures such as transformers. Vaswani et al. (2017) introduced the transformer model, which has become a cornerstone in natural language processing due to its ability to capture long-range dependencies through self-attention mechanisms. Recent research has continued to refine transformer architectures, with improvements in initialization playing a crucial role in achieving state-of-the-art performance (Liu et al., 2020).
In summary, our work contributes to the ongoing exploration of initialization methods by introducing a scheme that enhances the emergent properties of neural networks. By drawing on the principles of emergence and leveraging recent advancements in the field, we provide a novel perspective on how initialization can impact the learning dynamics and performance of deep models across various architectures and tasks.
@@ 3 Method
Emergence in machine learning refers to the phenomenon where complex, higher-level behaviors and structures arise from the interactions of simpler elements within a system. This concept, rooted in various scientific disciplines, has been increasingly recognized for its potential to enhance our understanding and development of neural networks. In our formulation, emergence is quantified through a graph-theoretical framework, which examines the network topology and the interactions between nodes to provide a clear and structural metric for emergent behavior.
A structural definition of emergence offers several key benefits. First, it allows for a more precise and quantifiable understanding of how complex behaviors develop in neural networks. This is particularly useful for designing and optimizing network architectures, as it provides clear metrics that can be targeted during the training and initialization phases. Second, a structural approach facilitates better theoretical insights, enabling researchers to draw more concrete connections between network topology and emergent properties. This can lead to more effective and efficient network designs, ultimately improving performance across various tasks.
Emergence in this context is inherently multiscale. It involves interactions across different scales of the network, where $G$ represents one scale and $H$ represents a higher scale. Emergence appears only when viewed from this multiscale perspective, as it captures the complexity arising from the network’s hierarchical structure.
@@
In our graph-theoretical framework, the emergence value $E$ of a neural network is defined based on the number of paths from nodes at scale $G$ to nodes at scale $H$. This definition captures the essence of multiscale interactions within the network. The more paths that exist between these scales, the greater the degree of emergence.
Mathematically, this can be expressed as:
$\text{Emergence}(G,H)$ $=\sum_{x\in G\setminus H}\#\text{paths in $G$ from $N_{H}(x)$ to $H$}.$ (1)
where:
@@
  • $H$ is the set of nodes at the higher scale,
This equation captures the emergent behavior by accounting for the connections and interactions between different scales within the network. The higher the value of $E$, the more interconnected the network is across scales, thereby increasing the likelihood of complex behaviors and traits emerging from the network.
Intuitively, a system with a higher value of $E$ has more extensive and interconnected pathways through which information can propagate across different scales. This interconnectedness facilitates the development of intricate patterns and features within the network, enabling it to capture and represent more complex relationships in the data. As a result, networks with higher emergence values are better equipped to learn and generalize from diverse and intricate datasets, leading to improved performance across various machine learning tasks.
Our approach leverages this definition to modify the initialization process of neural networks, aiming to enhance their emergent properties from the outset. By doing so, we can achieve significant improvements in network performance, as demonstrated in our experimental results.
@@
To adapt the definition of emergence in our context, we should give a clear definition of $G$ and $H$ in the mathematical definition above. Since emergence here evaluates the potential/ ability for emergent traits when we observe system $G$ from a higher level $H$, here we want $G$ to represent the model itself, and $H$ to be some certain features of the model. In the paper, we adopt the setting that $H$ is the nodes in $G$ that are still active in the training process, where the criteria for active nodes is the set of nodes whose activation is greater than a threshold set close to 00. This sorted out the nodes that are not actively participating in the computational process. The set of active nodes thus in a sense represent the learning task, thus we can tie emergence with the performance of the network in a learning process. This fits in our framework of emergence, where part of the system is being neglected after the learning process, thus the learning process represents the $\Phi$ where partial observation is carried out, and the properties of $H$ represents the emergent abilities of the network.
@@
Lemma: Emergence function $E(a_{1},\ldots,a_{N})$ increases when $a_{1}\ldots a_{i}$ gets smaller and $a_{i}\ldots a_{N}$ gets larger, where $i$ is the largest integer such that
$\sum_{k=1}^{N}a_{k}-a_{i}-\sum_{k=1}^{i-1}n_{i}>0.$ (4)
The proof is simple: when $a_{i}$ decrease by $1$, then the number of paths from layer $i$ to the dead nodes in latter layers will increase by $a_{i+1}+\cdots+a_{N}$, and the number of paths from earlier layers to layer $i$ will decrease by $(n_{1}-a_{1})+\cdots+(n_{i-1}-a_{i-1})$.
So the net increase of path will be
$\Delta$ $=a_{i+1}+\cdots+a_{N}-[(n_{1}-a_{1})+\cdots+(n_{i-1}-a_{i-1})]$ $=\sum_{k=1}^{N}a_{k}-a_{i}-\sum_{k=1}^{i-1}n_{i}$ (5)
For the usual neural network structures, where for each layer most nodes are active, and later layers are smaller, the balance usually occurs when $i\approx N/2$.
Let us consider now why the network with the configuration above has stronger emergence: for example, when the network is doing an image recognition task, the nodes in the later half of the layers are making important decisions on which category the image belongs to, so it needs to be more alive/ subject to more sensitive weight changes. However, for the nodes in the initial half of the layers, they are subject to greater weight changes, and they could be turned off to represent some global features of the image. Hence they could be more inactive/ subject to more drastic weight changes.
@@
  • Decrease the magnitude of weights in the first half layers by dividing a factor
  • Increase the magnitude of weights in the second half layers by multiplying a factor
@@
In our experiments, we see (Figure 1) that this initialization is indeed leading to a better performance. In particular, we show the correlation between emergence and performance. Based on our theory, we have an increase in emergence, even when only the magnitude of weights of the first half layers decreases, or the magnitude of weights of the second half increases.
We also note that, such choice of layer magnitudes is mimicking the "domino effect", as illustrated in the figure below. The increase of energy level for each piece can set off a cascade effect.
Figure 1: The increase of weight magnitude through the layers is mimicking how the energy level through the pieces of a domino is increasing.
Figure 2: Comparison of training loss and emergence of Xavier, Kaiming and our initialization schemes.
Emergence value: Xavier: $5.03e^{8}$, Kaiming: $5.99e^{8}$, Ours: $10.87e^{8}$

### v3

@@ Abstract
Emergence in machine learning refers to the spontaneous appearance of complex behaviors or capabilities that arise from the scale and structure of training data and model architectures, despite not being explicitly programmed. We introduce a novel yet straightforward neural network initialization scheme that aims at achieving greater potential for emergence. Measuring emergence as a kind of structural nonlinearity, our method adjusts the layer-wise weight scaling factors to achieve higher emergence values. This enhancement is easy to implement, requiring no additional optimization steps for initialization compared to GradInit. We evaluate our approach across various architectures, including MLP and convolutional architectures for image recognition and transformers for machine translation. We demonstrate substantial improvements in both model accuracy and training speed, with and without batch normalization. The simplicity, theoretical innovation, and demonstrable empirical advantages of our method make it a potent enhancement to neural network initialization practices. These results suggest a promising direction for leveraging emergence to improve neural network training methodologies. Code is available at: https://github.com/johnnyjingzeli/EmergenceInit.
@@ 1 Introduction
Emergence, in general, refers to the phenomenon where complex behaviors and properties arise from the interactions of simpler elements within a system. In machine learning, emergence has been studied as the nonlinear increase in system performance as the system’s size increases, exemplified by the emergent abilities of large language models. These emergent behaviors are crucial for enabling neural networks to perform complex tasks such as image recognition, natural language processing, and strategic game playing (Brown, 2020; Kaplan et al., 2020; Radford et al., 2019).
Although the concept of emergence has been observed in various fields and disciplines—such as phase transitions in physics and emergent structures and functions in biological networks—a unifying trait of these emergent phenomena is their association with nonlinearity. Generalizing the notion of a nonlinear function in calculus, this nonlinearity implies the disproportionate increase in system behavior when moving from the partial to the overall structure of the system.
A natural question regarding emergence is: what kind of system has a stronger potential for emergence? It is generally appealing to link the emergent function with the structure. To address this question, (Li et al., 2023) developed a measure of emergence based on network structure. This measure, which quantifies how much emergence a system can sustain, suggests a design principle for neural networks, enabling us to tune the network structure to maximize emergence.
Based on this measure, we propose a neural network initialization scheme that encourages emergence. The initialization of network parameters significantly impacts the training stability and performance of deep neural networks. Initializations that prevent gradient explosion or vanishing during backpropagation played a key role in the early successes of feed-forward networks (Glorot & Bengio, 2010; He et al., 2015). However, it remains theoretically challenging to link a network’s initialization with its training dynamics, especially for structure- and dataset-agnostic initialization schemes (Glorot & Bengio, 2010; He et al., 2015; Saxe et al., 2013; Mishkin & Matas, 2015; Zhu et al., 2021; Gilmer et al., 2021).
Our motivation differs from existing literature, which emphasizes network stability. By initializing networks with a stronger potential for emergence, we increase the likelihood of exhibiting emergent behaviors and patterns during training. This nonlinearity-based emergence suggests that the network structure and functionality are more susceptible to change, intuitively leading to larger training gradients.
We show that network architectures with stronger emergence, based on our measure, exhibit patterns of increasing activation, resembling natural emergent structures like dominos, where initial perturbations can lead to significant global changes, aligning with the general notion of emergence.
In this paper, we introduce a new initialization scheme for neural networks that leverages the concept of emergence. Our method adjusts layer-wise variance parameters to achieve higher emergence values compared to traditional methods like Xavier and Kaiming initialization (Glorot & Bengio, 2010; He et al., 2015). This approach is particularly appealing because it is straightforward to implement, requiring only minor modifications to existing initialization techniques without necessitating additional optimization or training steps.
@@ 2 Related Work
The initialization of neural networks has been a critical area of research, influencing the stability and speed of training, as well as the ultimate performance of the models. Traditional initialization schemes, such as Xavier (Glorot & Bengio, 2010) and Kaiming (He et al., 2015), have laid the foundation for effectively training deep networks by mitigating issues related to vanishing and exploding gradients. Xavier initialization aims to keep the scale of the gradients approximately the same in all layers, while Kaiming initialization, specifically designed for ReLU activations, helps to maintain the variance of activations throughout the layers. Both methods have proven to be fundamental in training deep networks but do not explicitly account for emergent properties within the networks.
@@
Recent research has explored more sophisticated initialization strategies that leverage the structural and statistical properties of neural networks. For instance, (Saxe et al., 2013) studied the dynamics of signal propagation in deep networks, highlighting the importance of properly scaling the initial weights to ensure efficient training. Additionally, (Mishkin & Matas, 2015) proposed a layer-sequential unit-variance (LSUV) initialization that iteratively adjusts the weights to achieve unit variance across all layers, further improving convergence.
@@
The concept of emergence, where complex behaviors arise from simple interactions within a system, has also been examined in the context of neural networks. Emergent properties have been shown to play a crucial role in the development of robust and adaptive models. Research by (Olah et al., 2020) illustrated how higher-level features and behaviors emerge in deep networks as a result of training on large datasets. This phenomenon underscores the potential for leveraging emergent properties to enhance network performance.
Despite these advancements, several challenges and limitations persist. Traditional initialization methods, while effective at preventing gradient-related issues, do not account for the complex emergent properties that can significantly influence network performance. More sophisticated methods, such as LSUV, improve convergence but may require iterative adjustments that complicate the initialization process.
The concept of emergence itself, although promising, is not yet fully understood or integrated into standard practices for network initialization. While studies like those by (Adam, 2017; Li et al., 2023) have made significant strides, there is still a need for practical methods that can harness emergent properties effectively from the outset of training.
@@
Furthermore, the role of initialization in specific architectures such as transformers remains an area of active research. (Vaswani, 2017) introduced the transformer model, which has become a cornerstone in natural language processing due to its ability to capture long-range dependencies through self-attention mechanisms. However, recent research continues to refine transformer architectures, with improvements in initialization playing a crucial role in achieving state-of-the-art performance [15] (Liu et al., 2020).
@@ 3 Method
Emergence fundamentally arises from the observation of a system from a higher scale. We build our definition of emergence on the notion of nonlinearity as the information passed to higher scales. Two key conceptual components are necessary to qualitatively describe emergent effects within the framework proposed by (Adam, 2017). The first is a notion of interaction or local computation among the components of a system. For example, the communication and propagation of information among nodes or subnetworks in the neural network. The second is the notion of interactional effects, which equips each system with an observable, for example, attaches network with its performance or abilities. These kinds of interactional effects are almost always associated with partial observations, or a simplification and integration of lower — more foundational or granular —levels or scales in the system that result in a ’loss of information’ or pattern/ feature formation at a higher level.
Figure 1: An illustration of emergence in the hierarchical system.
With these two ingredients, we can define emergence as a partial observation of interacting and interconnected components within a system that cannot be explained by known interactions that produce or result in partial observations of the components. This notion agrees with the intuitive understanding of emergence that some properties of the interconnected components cannot be decomposed or reduced to combinations of known properties of the constituent components, i.e. that the whole is more than the sum of its parts. This notion of emergence is the foundation on which our work in this paper, building on the framework first proposed in (Adam, 2017), develops a mathematical definition and computational measure of emergence.
To formalize these ideas, we begin by representing the interactions between components as an operation $\vee$, where $s_{1}\vee s_{2}$ represents a new interconnected system of subsystems $s_{1}$ and $s_{2}$. Interactional effects are described by the mapping $\Phi$ that sends a system to its partial observation or interactional effect at a higher scale, in some cases corresponding to a coarse graining scheme (Rosas et al., 2024). Emergent effects are sustained whenever the observation of the combined system cannot be explained by the observation of the separate components. Mathematically,
Definition A system sustains emergent effects when the following inequality is satisfied:
$\Phi(s_{1}\vee s_{2})\neq\Phi(s_{1})\vee\Phi(s_{2}),$ (1)
for some constituent subsystems $s_{1}$ and $s_{2}$.
This definition essentially captures emergence as a kind of "structural nonlinearity". Let’s consider the simple case when $\Phi$ is simply a smooth function $f:\mathbb{R}\rightarrow\mathbb{R}$ and the interaction $\vee$ is simply taking the average, $s_{1}\vee s_{2}=(s_{1}+s_{2})/2$. Then we realize that the extent to each $\Phi(s_{1}\vee s_{2})$ differs from $\Phi(s_{1})\vee\Phi(s_{2})$ is just $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$, which is related to how nonlinear the function $f$ is, and can be studied by the derivatives of $f$, in particular, the second order derivative, since $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$ can be approximated by $\frac{|s_{2}-s_{1}|^{2}}{4}|f^{\prime\prime}(\xi)|$ for some $\xi$ between $s_{1}$ and $s_{2}$. Now when $\Phi$ is a functor [26], which captures the cross-scale information flow in real world systems, we want an analogue to derivatives to apply this idea, and this naturally leads to the concept of a derived functor in homological algebra[26]. We can also see that Definition captures the structural nonlinearity in emergence, the nonlinearity of system’s behavior and functionality as the system’s structure changes, or as we go from components to parts of the system to the whole system. This is a general mathematical definition of emergence, as first given in (Adam, 2017)  [^1] In (Adam, 2017) the term ”generativity” is used instead of emergence. These two terms are often considered interchangeable.. Note that when studying the emergence of a specific system, the interaction $\vee$ and interactional effect $\Phi$ need to be chosen carefully.
Examples of emergence in machine learning Emergence or generativity has been a rising concept in machine learning, for example, (Wei et al., 2022a, b; Du et al., 2024). Emergent abilities of large language models, for example, (Wei et al., 2022a), commonly conceived as the new properties/ abilities of the larger models that do not exist in smaller models. If we consider $s_{1}$ and $s_{2}$ as two smaller models, $s_{1}\vee s_{2}$ as combining two smaller models into a larger model by, for example, techniques in ensemble learning (Mohammed & Kora, 2023), and $\Phi$ as the mapping that reflects the properties/ abilities of the model, that is, $\Phi(s)$ is the ability acquired by the model $s$. Then $\Phi(s_{1}\vee s_{2})$ is the properties/ abilities of the combined model and $\Phi(s_{1})\vee\Phi(s_{2})$ can be interpreted as a summation of the properties/ abilities of each small model. Then the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ can reflect the emergent properties/ abilities that result in the nonlinear increase of performance, related to the performance in (Wei et al., 2022a). The difference can also be related to generalizability, where $s_{1}$ and $s_{2}$ are two data sets, when the model trains on two data sets, it is usually different from training the model on separate datasets.
Based the mathematical theory of emergence in (Adam, 2017), the structural difference in (1) between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$, can be evaluated through computing the mathematical structure of derived functor $R^{1}\Phi$, see (Adam, 2017; Rotman & Rotman, 2009). (Li et al., 2023) gives the following result that computes $R^{1}\Phi$, generalization of derivative, where the input is the mathematical structure of quiver representation.
Theorem (Proposition 5.3 in (Li et al., 2023)) Given the functor $\Phi$ which preserves partial structure in a quiver representation $W$ by deleting a set of edges $E$, the derived functor of $\Phi$ is
$R^{1}\Phi(W)=\bigoplus_{a\in E}\Phi(W(ta)\otimes P_{ha})$ (2)
where $ta$ is the tail of edge $a$ (the starting node), $ha$ is the head of edge $a$ (the ending node), $W(ta)$ is the vector space associated to node $ta$, $P_{ha}$ is the vector space spanned by all paths originating from node $ha$.
Proof.
The proof of this theorem is given in the appendix. ∎
This theorem computes $R^{1}\Phi$, which evaluates the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$, thus encodes the potential of a system for emergence. Taking advantage of this theorem, we can take the dimension of $R^{1}\Phi(W)$ as a numerical approximation of the potential for emergence of $W$ when the network interact with other networks. Given a network $G$, and a sub-network $H$ which represents its effect or observation under $\Phi$, where their relation are shown as follows:
[…]
then we have the following measure of emergence for networks:
$\text{Emergence}(G,H)$ $=\sum_{x\in G\setminus H}\#\text{paths in $H$ from $N_{H}(x)$ to $H$},$ (3)
where $H$ represents the part of network structure being preserved by $\Phi$, the partial observation. $N_{H}(x)$ is the set of downstream neighbors of $x$ in $H$.
@@
Emergence in this context is inherently multiscale. It involves interactions across different scales of the network, where $G$ represents one scale and $H$ represents a higher scale. Emergence appears only when viewed from this multiscale perspective, as it captures the complexity arising from the network’s hierarchical structure. In our graph-theoretical framework, the emergence value $E$ of a neural network is defined based on the number of paths from nodes at scale $G$ to nodes at scale $H$. This definition captures the essence of multiscale interactions within the network. The more paths that exist between these scales, the greater the degree of emergence.
@@
  • $H$ is the set of nodes at the higher scale.
Our measure captures the emergent behavior by accounting for the connections and interactions between different scales within the network. The higher the value of $E$, the more interconnected the network is across scales, thereby increasing the likelihood of complex behaviors and traits emerging from the network. Intuitively, a system with a higher value of $E$ has more extensive and interconnected pathways through which information can propagate across different scales. This interconnectedness facilitates the development of intricate patterns and features within the network, enabling it to capture and represent more complex relationships in the data. As a result, networks with higher emergence values are better equipped to learn and generalize from diverse and intricate datasets, leading to improved performance across various machine learning tasks.
Our approach leverages this definition to modify the initialization process of neural networks, aiming to enhance their emergent properties from the outset. By doing so, we achieved significant improvements in network performance, as demonstrated in our experimental results.
@@
In machine learning setting, one modeling approach is to consider $\Phi$ as the training process, since emergence here evaluates the potential/ ability for emergent traits when we observe system $G$ from a higher level $H$, here we want $G$ to represent the model itself, and $H$ to be some certain features of the model. In the paper, we adopt the setting that $H$ is the nodes in $G$ that are still active in the training process, where the criteria for active nodes is the set of nodes whose average activation on all input data is greater than a threshold. This sorted out the nodes that are not actively participating in the computational process/ representing features. The set of active nodes thus in a sense represent the learning task, thus we can tie emergence with the performance of the network in a learning process. This fits in our framework of emergence, where part of the system is being neglected after the learning process, thus the learning process represents the $\Phi$ where partial observation is carried out, and the properties of $H$ represents the emergent abilities of the network.
@@
Lemma: Emergence function $E(a_{1},\ldots,a_{N})$ increases when $a_{1}\ldots a_{i}$ gets smaller and $a_{i+1}\ldots a_{N}$ gets larger, where $i$ is the largest integer such that
$-n_{i-1}+n_{i+1}+n_{i+1}n_{i+2}+n_{i+1}n_{i+2}n_{i+3}+\cdots+n_{i+1}n_{i+2}\cdots n_{N-1}n_{N}>0.$ (6)
Proof: Consider the case where all the layers before $i_{th}$ layer are fully inactive and all the layers after $i_{t}h$ layer are fully active, in other words, $a_{k}=0$ for $k<i$ and $a_{k}=n_{k}$ for $k>i$. Then when $a_{i}$ decrease by $1$, the number of paths from previous layers to layer $i$ will decrease by $n_{i-1}$, the number of paths from layer $i$ to latter layers will increase by $n_{i+1}+n_{i+1}n_{i+2}+n_{i+1}n_{i+2}n_{i+3}+\cdots+n_{i+1}n_{i+2}\cdots n_{N-1}n_{N}$. So the net increase of paths will be
$\Delta$ $=-n_{i-1}+n_{i+1}+n_{i+1}n_{i+2}+n_{i+1}n_{i+2}n_{i+3}+\cdots+n_{i+1}n_{i+2}\cdots n_{N-1}n_{N}.$ (7)
In a wide range of image recognition tasks, we propose to choose $i\approx N/2$ as it works for a wide range of neural network architecture blocks.
Let us consider now why the network with the configuration above has stronger emergence: for example, when the network is doing an image recognition task, the nodes in the later half of the layers are making important decisions on which category the image belongs to, so it needs to be more active/ subject to more sensitive weight changes. However, for the nodes in the initial half of the layers, they are subject to greater weight changes, and they could be turned off to represent some global features of the image. Hence they could be more inactive/ subject to more drastic weight changes.
@@
  • Decrease the magnitude of weights in the first half layers by dividing a factor $\alpha$
  • Increase the magnitude of weights in the second half layers by multiplying a factor $\alpha$
@@
In our experiments, we see in Figure 3 that this initialization is indeed leading to a better performance. In particular, we show the correlation between emergence and performance. Based on our theory, we have an increase in emergence, even when only the magnitude of weights of the first half layers decreases, or the magnitude of weights of the second half increases.
We also note that, such choice of layer magnitudes is mimicking the "domino effect", as illustrated in the figure below. The increase of energy level for each piece can set off a cascade effect. When the perturbation to input is small, like a Gaussian noise, the difference won’t propagate to the later layers in the network, but when the input change is significant, the difference will then propagate to the later layers, and be magnified. Hence network with strong emergence like this, will be robust against random noise while sensitive to feature change, suggesting a higher generalizability.
Figure 2: (A) The increase of weight magnitude through the layers mimicks how the energy level through the pieces of a domino is increasing. (B) Comparison of training loss and emergence of Xavier, Kaiming and our initialization schemes. Emergence measure: Xavier: $5.03e^{8}$, Kaiming: $5.99e^{8}$, Ours: $10.87e^{8}$